# DeToNATION: Decoupled Torch Network-Aware Training on Interlinked Online Nodes

### v3

@@ Results
Figure 2(a) shows the validation loss. Training loss is available in Appendix C. Random replication performs best, with compressions $\frac{1}{2}$ and $\frac{1}{4}$ being best, respectively. Subsequently, DeMo $\frac{1}{8}$ comes followed by DeMo $\frac{1}{4}$, $\frac{1}{16}$ and $\frac{1}{32}$. DiLoCo and Striding replication schemes generally convergences much slower. We see similar results for Random $\frac{1}{16}$ and $\frac{1}{32}$, where the signal seems to become too sparse, yet still learning.
Figure 2: Epoch train and validation loss for T5-Large on the Opus Books En-Fr subset.
(a) T5-Large Validation loss on the Opus Books En-Fr subset. Random and DeMo replication demonstrates strong performance.
(b) ViT-B Validation loss on Cifar100. DeMo and DiLoCo performs the best and very comparable.
@@ Discussion
We have shown that DeMo-SGD with Random-replication schemes is best for T5 of the replication schemes, with compression rates $\frac{1}{2}$ and $\frac{1}{4}$, whereas DeMo lands at second place, with compression rates $\frac{1}{8}$ and $\frac{1}{4}$, respectively. While DeMo enables more compression in its best-performing configurations, Random yields lower loss and is faster in practice. Striding and DiLoCo are converging slowly, making them less competitive here.
@@
We train a ViT-B (Patch $16$, 224224x224224) model for vision classification on the Cifar100 dataset (Krizhevsky, Hinton et al. 2009), for 20 epochs, in a Hybrid-FSDP setup with $2$ nodes with $8$ accelerators each. experiments are run with DeMo-SDG as the underlying optimizer with a learning rate of $1e^{-5}$.
@@ Results
We report our findings in Figure 2(b). Training loss is available in the Appendix D. DeMo replication with a compression rate of $\frac{1}{2}$ and $\frac{1}{4}$ demonstrates the highest performance (comparable), closely followed by DeMo at $\frac{1}{16}$, with DiLoCo $\frac{1}{2}$ in-between the two. We generally see that the Random scheme struggles quite clearly in this domain. Striding replication even performs better than Random.
@@ Discussion
Contrary to the encoder-decoder architecture on a next-token-prediction training with superior Random replication scheme, DeMo replication is superior here. We hypothesis that this is because fast moving momenta is more suited for this task, as less pronounced features doesn’t contribute to the learning-objective, potentially inducing noise, whereas these small nuances in language modeling is very valuable. Randomly chosen indices can include feature-information not important for the implicit supervised contrastive learning. DiLoCo with a quite low compression follows DeMo’s performance closely, while taking a hit on the performance when increasing the compression rate. Striding does follow DiLoCo’s performance closely, we hypothesis that the structure provided by this scheme work in highly structured data, that images are.
@@
We train the OLMo2 decoder-only model for causal language modeling, in a Hybrid-FSDP setup with $2$ nodes with $4$ accelerators each for 1B parameters. We train on the Dolma v1.6 dataset (Soldaini et al. 2024) for 10K steps. We use the OLMo’s standard parameters [^1] https://github.com/allenai/OLMo/blob/main/configs/official-0425/OLMo2-1B-stage1.yaml, with the chance of 4% warm-up steps. Each node is composed by a custom Ampere A100 64GB GPU, with 2x dual-port HDR network interface (400Gbps aggregated). The nodes are connected in a infiniband-based Dragonfly+ topology with a low latency interconnect with 200Gb/s. The purpose of our experiments is to demonstrate FlexDeMo’s reduction in communication overhead in comparison with and Hybrid-FSDP and a conventional AdamW optimizer (Kingma and Ba 2015; Loshchilov and Hutter 2019).
@@ Results
We report our findings in Figures 3 and 4. All of the FlexDeMo with DeMo replication demonstrates the best performance, with interestingly the compression-rate $\frac{1}{32}$, being the best, closely followed by $\frac{1}{16}$, demonstrating some interesting artifacts of high compression. Random-replication is next-best, and better than DeMo with compression $\frac{1}{4}$, which performs comparable to Random $\frac{1}{16}$. Hybrid-FSDP with conventional AdamW (red), performs better all DiLoCo configurations, but worse than all DeMo and Random configurations. Strided replication generally yields unstable training and uncompetitive results. In Figure 4, we see that all replicator-types across all compression-rates is significantly faster than conventional AdamW, which employs full replication. Conducting timings are quite unreliable on a HPC, as it depends on the network-congestion during the time of the experiment, preventing us to make exact comparisons between the replicators.
Figure 3: OLMo2 1B train loss (zoomed) over 10K training steps on Dolma v1.6 using different replicators and compression rates. All experiments, except the Hybrid-FSDP baseline with AdamW on 2 nodes, use DeMo-SGD.
@@ Discussion
Similarity, as with the ViT-archtiecture the DeMo-replicator performs best for OLMo-architecture with compression rates $\frac{1}{32}$ and $\frac{1}{16}$, being the best. This is somewhat closely followed the Random replication scheme, however with larger compressions on $\frac{1}{4},\frac{1}{8}$ and $\frac{1}{2}$, respectively, making them even less attractive to the DeMo variant. Interstingly, we do not see an optimal replication-scheme within each We demonstrated the impact of FlexDeMo with all the replicators compared to conventional Hybrid-FSDP and AdamW across only 2 nodes on a HPC. We see that all the replicator schemes yields training-times around approx. 2.6 times faster than the Hybrid-FSDP with AdamW setting. As timing on HPC systems is dependent on the current network congestion, we refrain from commenting on differences between the replicators, as they probably will alternate when reproducing. However, this does not make the big relative margin to the conventional setup uncertain and paves the way for better utilizing GPU allocations and lowering the environmental impact significantly. Thus, with the hybrid-sharding strategy, the all_gather is done only once per node, instead of once per accelerator in the original decoupled momentum implementation, which is clearly beneficial. With introducing hybrid-sharding, we need to use the gather-operation intra-node, which is reasonably fast, but can be optimized by implementing hooks and streams, which we leave for future work.
@@
In this experiment we investigate the performance and efficiency of FlexDeMo at scale. Our setup is based the experiments in Figure 3, employing the Demo-SGD optimizer, and both DeMo and Random-replication schemes with a $\frac{1}{32}$ compression rate and a learning-rate of $1e^{-3}$.
@@ Results
We report our findings in Figure 5. Similarly, we see that Random replication performs worse, than DeMo replication, which performs almost comparable to conventional AdamW (full-sync). However, DeMo is significantly slower. We attribute the to DeMo’s dependence on the all_gather operation, which prevents is to scale meaningfully, shown In Figure 6. The values on the $y$-axis is smaller than in 3, due to only tracking on rank 0, to not induce a bottleneck for tracking.
@@ Discussion
In Figures 5 and 6, we investigate the scale both Random and DeMo replication schemes, powered by the DeMo-SGD optimizer and comparing against conventional Hybrid-FSDP with AdamW optimization. While DeMo replications performance is clearly best, there is a big drawback in the time-scaling of the scheme, due to the all_gather operation not scaling, as one would expect, due to the blocking nature of the operation. However, Random is around 64% faster than the conventional setup. This could call for 2-stage distributed training setup, employing Random replication for the majority of the training, employing the conventional setup for a subsequent stage.
Figure 5: OLMo2 1B train loss (zoomed) over 10K training steps on Dolma v1.6 using DeMo-SGD with a 1/32 compression rate on 64 nodes. We compare repl. schemes DeMo and Random against the Hybrid-FSDP + AdamW baseline.
@@
General Discussion and Conclusion
Through a set of experiments, we have validated important hyper-parameters, investigating the characteristics of FlexDemo, with DeMo replication, including communication data type, underlying optimizer, the question on the sign function and TopKK. We do expect parameters to be architecture and potentially domain-dependent.
We demonstrate that using sign before synchronizing is a corner-stone in this optimization scheme. Immediately, this tells us that direction is more important than magnitude for our case of decoupled optimization. This has also been studied in prior work (Bernstein et al. 2018). The ternary system opens up for the possibility to compress the data even more, assuming it can be done fast enough.
In experiments with full control over the network bandwidth between the nodes, we have shown how Random replication scales across different replicators and compression rates . This, with the DeMo-SGD optimizer, is not surprising, as SGD employs fewer parameters per model parameter than AdamW.
We have introduced a decoupled variant of AdamW, investigating a version not sharing the first and second moments. However, we demonstrate in Figure 1, that Decoupled-AdamW is not generally superior and only outperforms DeMo-SGD when compared on the DiLoCo-scheme. We did not develop this optimizer further, as we in OLMo2 demonstrated superior performance relative to conventional Hybrid-FSDP with AdamW.
We have studied the Striding and Random replication schemes we introduced to facilitate FlexDeMo and challenged experimental choices made in DeMo. Using any of these two schemes enables us to utilize more of the bandwidth sending actual data and not the corresponding indices. We demonstrate that DeMo works best on the ViT and decoder-architectures, whereas Random performs best in the encoder-decoders architecture calling for investigation into more architecture-specific replications.
Throughout our experiments we have experimented with both NCCL and RCCL communications-backends, demonstrating DeToNATION and FlexDeMo across both platforms. Overall, this opens an avenue of future research of decoupled training-techniques, extending these schemes but also, importantly, investigating sparse optimization in different domains. This is important, as we demonstrated that different replicators are optimal for different tasks. For future work, we leave an open question of weather the type of replicator depends on the task or architecture.
@@
Introducing FSDP requires the gathering of full parameters, hence an initial reduce_scatter-operation on each gradient. PyTorch’s FSDP implementation does this when the gradients become available during the backward-pass, which we have disables with no_sync. However, this can be implemented using hooks, which will enable starting already during the backward-pass. Further, we believe that improving communication with CUDA streams in a general asynchronous synchronization scheme contains a substantial efficiency potential. Introducing FSDP2222https://docs.pytorch.org/docs/stable/
distributed.fsdp.fully_shard.html#pytorch-fsdp2-fully-shard
or SimpleFSDP (Zhang et al. 2024) will also speed up the basic communication operations. These implementations will make the groundwork relaxing the memory-requirements even further, more comfortably enabling cross-node sharding and, thus, supporting very large models. Further, this opens for the feasibility of connecting one or more HPC systems when training, simplifying collaborations and utilizing the existing resources.
@@ Conclusion
In conclusion, we have shown that FlexDeMo successfully extends DeMo optimization to the setting of hybrid sharded FSDP. Our experiments indicate results better than AdamW, while at the same time being significantly faster, demonstrating the potential to reduce the bandwidth needed for communication between nodes. We can conclude that DeMo-SGD proves to be the superior underlying optimizer in most setups, while the underlying replicator is more dependent on given architecture and task. We also introduced Decoupled AdamW. DeMo replication scales significantly worse than the conventional scheme, whereas Random, performing a little worse, keeps delivering across many nodes. Overall, this enables FlexDeMo to be used to optimize LLMs that do not fit within the memory of one accelerator and further supports training LLMs in low-resource network settings, ultimately lowering the entry point for both practitioners and researchers to participate in the development of future models and use cases, while offering more compute from the same GPU hours, potentially reducing the overall environmental impact.
Appendix A A: Communication DeMo vs. FlexDeMo

### v4

@@ Results
Figure 1 shows the validation loss. Training loss is available in Appendix C. Random replication performs best, with compressions $\frac{1}{2}$ and $\frac{1}{4}$ being best, respectively. DeMo $\frac{1}{8}$ comes in second followed by DeMo $\frac{1}{4}$, $\frac{1}{16}$ and $\frac{1}{32}$. DiLoCo and Striding replication schemes generally converge much slower. We see similar results for Random $\frac{1}{16}$ and $\frac{1}{32}$, where the signal seems to become too sparse.
Figure 1: T5-Large Validation loss on the Opus Books En-Fr subset. Random and DeMo replication demonstrates strong performance.
Figure 2: ViT-B Validation loss on Cifar100. DeMo and DiLoCo perform best and similar to each other.
@@ Discussion
We have shown that DeMo-SGD with Random-replication schemes is best for T5 of the replication schemes, with compression rates $\frac{1}{2}$ and $\frac{1}{4}$, whereas DeMo takes second place, with compression rates $\frac{1}{8}$ and $\frac{1}{4}$, respectively. While DeMo enables more compression in its best-performing configurations, Random yields lower loss and is faster in practice. Striding and DiLoCo are converging slowly, making them less competitive here.
@@
We train a ViT-B (Patch $16$, 224224x224224) model for vision classification on the Cifar100 dataset (krizhevsky2009learning). Experiments are run with DeMo-SGD as the underlying optimizer with a learning rate of $10^{-5}$.
@@ Results
We report our findings in Figure 2. Training loss is available in the Appendix D. DeMo replication with compression rates $\frac{1}{2}$ and $\frac{1}{4}$ demonstrates the highest performance (comparable), closely followed by DeMo at $\frac{1}{16}$, with DiLoCo $\frac{1}{2}$ in-between the two. We generally see that the Random scheme struggles in this domain. Even Striding replication performs better than Random.
@@ Discussion
Contrary to the encoder-decoder architecture on a next-token-prediction training with superior Random replication scheme, DeMo replication is superior here. We hypothesis that this is because fast-moving momenta is more suited for this task, as less pronounced features do not contribute to the learning-objective, potentially inducing noise, whereas these small nuances seem to be valuable for language modeling. Randomly chosen indices can include feature-information not important for the implicit supervised contrastive learning. DiLoCo with a low compression follows DeMo’s performance closely, while taking a hit on the performance when increasing the compression rate. Striding does follow DiLoCo’s performance closely, we hypothesis that the structure provided by this scheme work in highly structured data, that images are.
@@
We train the OLMo2 decoder-only model for causal language modeling, in a Hybrid-FSDP setup with $2$ nodes with $4$ accelerators each for 1B parameters. We train on the Dolma v1.6 dataset (dolma) for 10K steps. We use the OLMo’s standard parameters [^1] https://github.com/allenai/OLMo/blob/main/configs/official-0425/OLMo2-1B-stage1.yaml, with the chance of 4% warm-up steps. Each node is composed by a custom Ampere A100 64GB GPU, with 2x dual-port HDR network interface (400Gbps aggregated). The nodes are connected in a InfiniBand-based Dragonfly+ topology with a low latency interconnect with 200Gb/s.
@@ Results
We report our findings in Figures 3 and 4. All of the FlexDeMo with DeMo replication demonstrates the best performance, with interestingly the compression rate $\frac{1}{32}$ performing best, closely followed by $\frac{1}{16}$, displaying some interesting artifacts of high compression. Random-replication is second-best, and better than DeMo with compression $\frac{1}{4}$, which is comparable to Random $\frac{1}{16}$. Hybrid-FSDP with conventional AdamW (red), performs better than all DiLoCo configurations, but worse than all DeMo and Random configurations. Strided replication generally yields unstable training and uncompetitive results. In Figure 4, we see that all replicator-types across all compression rates are substantially faster than conventional AdamW, which employs full replication. Conducting wall-time measurements on an HPC is not perfectly reliable, as they can be affected by the network-congestion during the time of the experiment, preventing exact comparisons between replicators. However, the measurements still provide insights in the general tendencies, and scale of expected runtime.
Figure 3: OLMo2 1B train loss (zoomed) over 10K training steps on Dolma v1.6 using different replicators and compression rates. All experiments, except the Hybrid-FSDP baseline with AdamW on two nodes, use DeMo-SGD.
@@ Discussion
Similarly as for the ViT architecture the DeMo replicator performs best for OLMo architecture with compression rates $\frac{1}{32}$ and $\frac{1}{16}$, being the best. This is somewhat closely followed the Random replication scheme, however with larger compressions on $\frac{1}{4},\frac{1}{8}$ and $\frac{1}{2}$, respectively, making them even less attractive to the DeMo variant. Interstingly, we do not see an optimal replication scheme within each We compared all FlexDeMo replicators to Hybrid-FSDP with AdamW on two HPC nodes. We see that all the replicator schemes yields training times around 2.6 times faster than the Hybrid-FSDP with AdamW setting. As timing on HPC systems is dependent on the current network congestion, we do not comment on smaller differences between the replicators. However, this does not make the large relative margin to the conventional setup uncertain, and paves the way for better utilizing GPU allocations and lowering the environmental impact substantially. We can thus see that it is clearly beneficial that the all_gather is done only once per node with the hybrid-sharding strategy, instead of once per accelerator in the original decoupled momentum implementation. With hybrid-sharding, we still need to use the gather-operation intra-node, which is reasonably fast, but can be optimized by implementing hooks and streams. We leave this for future work.
Replicator bandwidth usage
Setup
Here we measure the actual bandwidth usage for the DeMo replicator, Random replicator and full replication, while training a T5-small model with the French-English subset of the OpusBooks dataset for 20 epochs. The setup is based on 8 RTX A6000 GPUs, in two nodes of 4 cards each, connected together with a 10 Gbps network interface. The environment is entirely controlled, with no other users, to allow truthful bandwidth measurements. The replicators are running at a compression rate of $1/16$
Results and Discussion
We measured an average bandwidth usage of $1070$ Mbps for full replication, $291$ Mbps for DeMo, and $152$ Mbps for Random. That is, DeMo uses about twice the bandwidth as Random, and full replication uses 7 times more. Full replication uses close to the absolute limit of the cable, and we expect that it is indeed bottlenecked by the bandwidth, explaining why it doesn’t reach 16 times the use of the other replicators. We expect that DeMo use twice the bandwidth of Random, as DeMo is sharing both the gradient values, and the index of the shared gradients, due to the nature of the compression scheme. Random avoids this, by seeding the random generator on all nodes, so the indices are calculated locally, and thus not transferred. Consequently, Random can effectively share twice the amount of gradients for the same bandwidth cost.
Performance in congested networks
Setup
We use a 2 node setup, each with 4 RTX A6000 GPUs, and connected with 10 Gbps ethernet, to train T5-Large and ViT-B at different bandwidth limits. As most HPCs are subject to congested networks when training at scale, we simulate such a limited network, by limiting the bandwidth of the inter-node connection. Here, we impose the limit using Linux Traffic Control on the nodes. We analyze the average time per training step for $10$, $100$, 1,0001{,}000 and 10,00010{,}000 Mbps connections, and compare DeMo and Random replication, with underlying DeMo-SGD optimizer, with compression rates $\frac{1}{16}$ and $\frac{1}{32}$, against the Decoupled-AdamW optimizer (ours).
Results
We report our results for ViT in Figure 5, and for T5 in Appendix D. As expected, as the bandwidth decreases, the importance of the compression rate becomes more pronounced, where the compression of $\frac{1}{32}$ clearly demonstrates to be fastest followed by $\frac{1}{16}$ with $\frac{1}{1}$, being last, dictated by the amount of data to communicate, with Random DeMo-SGD being the fastest, as shown both in Figure 18 and 5. This is crucial at bandwidths lower than 500Mbps, a level of bandwidths comparable to those available in HPCs in practice. Random replication is approximately $3.33$ times faster than DeMo-replication at 1010Mbps, while being approx. $18$ times faster than Decoupled-AdamW with full replication (Figure 5). Generally we see that Random replication with compression $\frac{1}{16}$ scales similarly to DeMo replication with compression $\frac{1}{32}$, demonstrating exactly that Random replication is around twice as fast as DeMo, as expected from DeMo transferring twice the amount of data, at the same compression rate.
Figure 5: Average time per optimizer step for ViT-B on two nodes across employing different replicators using DeMo-SGD and Decoupled AdamW as base optimizers
@@ Results
In this experiment we investigate the performance and efficiency of FlexDeMo at scale. Our setup is based on the same experiments as the previous OLMo2 experiment, using the Demo-SGD optimizer, and both DeMo and Random-replication schemes with a $\frac{1}{32}$ compression rate and a learning rate of $10^{-3}$. We scale to 64 nodes à 4 GPUs.
@@
We report our findings in Figure 6. We see that Random replication performs worse than DeMo replication, which performs almost comparable to conventional AdamW (full-sync). However, DeMo is substantially slower. We attribute this to DeMo’s dependence on the all_gather operation, which prevents it to scale as favorable as the Random replicator. This is shown in Figure 7.
@@ Discussion
In Figures 6 and 7, we investigate the scale both Random and DeMo replication schemes, powered by the DeMo-SGD optimizer and comparing against conventional Hybrid-FSDP with AdamW optimization. While DeMo replications performance is clearly best, there is a big drawback in the time-scaling of the scheme, due to the all_gather operation scaling badly, as one would expect, due to the blocking nature of the operation. However, the Random replicator is around 64% faster than the conventional setup. This could call for 2-stage distributed training setup, employing Random replication for the majority of the training, using the conventional setup for a subsequent stage.
Figure 6: OLMo2 1B train loss (zoomed) over 10K training steps on Dolma v1.6 using DeMo-SGD with a 1/32 compression rate on 64 nodes. We compare replication schemes DeMo and Random against the Hybrid-FSDP + AdamW baseline.
@@
General Discussion
Through a set of experiments, we have validated important hyper-parameters, investigating the characteristics of FlexDemo, with DeMo replication, including communication data type, underlying optimizer, and choice of the sign function and TopKK. We do expect parameters to be architecture and potentially domain-dependent.
We demonstrate that using sign before synchronizing is a corner-stone in this optimization scheme. Immediately, this tells us that direction is more important than magnitude for our case of decoupled optimization. This has also been studied in prior work (bernstein2018signsgd). The ternary system enables compressing the data even more.
In experiments with full control over the network bandwidth between the nodes, we have shown how Random replication scales better across different replicators and compression rates. This, with the DeMo-SGD optimizer, is not surprising, as SGD employs fewer parameters per model parameter than AdamW.
We have introduced a decoupled variant of AdamW, investigating a version not sharing the first and second moments. However, we demonstrate in Figure 19 of the Appendix, that Decoupled-AdamW is not generally superior and only outperforms DeMo-SGD when compared on the DiLoCo replicator. We did not develop this optimizer further, as the OLMo2 experiment has shown superior performance of Hybrid-FSDP with AdamW.
We have studied the Striding and Random replication schemes we introduced to facilitate FlexDeMo and challenged experimental choices made in DeMo. Using any of these two schemes enables us to utilize more of the bandwidth sending actual data and not the corresponding indices. We demonstrate that DeMo works best on the ViT and decoder-architectures, whereas Random performs best in the encoder-decoders architecture.
Throughout our experiments, we have experimented with both NCCL and RCCL communications-backends, demonstrating DeToNATION and FlexDeMo across both platforms. Overall, this opens an avenue of future research of decoupled training-techniques, extending these schemes but also, importantly, investigating sparse optimization in different domains. This is important, as we demonstrated that different replicators are optimal for different tasks. Future work may study whether the ideal type of replicator depends on the task or architecture, potentially warranting the development of architecture-optimized replicators.
@@
Introducing FSDP requires an initial reduce_scatter operation on each gradient. PyTorch’s FSDP implementation performs this during the backward pass, which we disabled with no_sync. This can be implemented using hooks to enable overlap with the backward pass. Additionally, asynchronous communication with CUDA streams contains substantial efficiency potential. Adopting FSDP2 or SimpleFSDP (zhang2024simplefsdp) will further accelerate basic communication operations. These improvements will enable cross-node sharding for very large models and facilitate connecting multiple HPC systems for collaborative training.
@@ Conclusion
We have shown that FlexDeMo successfully extends DeMo optimization to hybrid sharded FSDP, achieving comparable or better results than AdamW while being significantly faster. DeMo-SGD has displayed superior performance as the underlying optimizer in most settings, while replicator choice depends on architecture and task. DeMo replication demonstrates the best convergence but scales poorly with node count due to all_gather overhead, whereas Random replication maintains efficiency at scale despite slightly worse loss. FlexDeMo enables training LLMs that exceed single-accelerator memory and supports low-bandwidth network settings, lowering barriers for practitioners and researchers while improving compute efficiency and reducing environmental impact.
Acknowledgments
We acknowledge EuroHPC JU for awarding the projects EUHPC_A04_086 access to Leonardo BOOSTER and DeiC-SDU-S5-202500013 for access to LUMI and Ordbogen A/S for providing resources for bandwidth constrained experiments. This research was in parts supported by the Danish Foundation Models (DFM) project.
Appendix A Reproducibility
[…]
Appendix B Reproducibility Checklist
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced: Yes
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results: Yes
  • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper: Yes
  • Does this paper make theoretical contributions? No
  • Does this paper rely on one or more datasets?: Yes
    • A motivation is given for why the experiments are conducted on the selected datasets Yes
    • All novel datasets introduced in this paper are included in a data appendix. No
    • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. Yes
    • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations Yes.
    • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. Yes
    • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. NA.
  • Does this paper include computational experiments? Yes.
    • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting Yes
    • Any code required for pre-processing data is included in the appendix. Yes.
    • All source code required for conducting and analyzing the experiments is included in a code appendix.Yes
    • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes Yes
    • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from No
    • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. Yes
    • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks Partial
    • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. No
    • This paper states the number of algorithm runs used to compute each reported result. Yes
    • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. No
    • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). No
    • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments Yes
Appendix C A: Communication DeMo vs. FlexDeMo